# Improving propulsive efficiency using bio-inspired intermittent locomotion

### v1

@@ Abstract
Many swimmers, especially small to medium-sized animals, use intermittent locomotion that differs from continuous swimming of large species. This type of locomotion, called burst-and-coast, is often associated with an energetic advantage. In this work, we investigate the intermittent locomotion inspired by fish locomotion but applied to a propeller. The energy consumption of burst-and-coast cycles is measured and compared to the continuous rotation regime. We show that a substantial drag contrast between the active and passive phases of the motion, as observed in fish, is critical for energy savings. Such a contrast can be obtained using a folding propeller that passively opens and closes as the propeller starts and stops rotating. For this reconfigurable propeller, intermittent propulsion is found to be energetically advantageous, saving up to 24% of the energy required to cruise at a given speed. Using an analytical model, we show that intermittent motion is more efficient than continuous motion when the drag reduction in the coast phase exceeds 65%. For fish-like locomotion, this threshold seems to be closer to 30%. A formal analogy allows us to explain the difference between propeller propulsion and fish locomotion.
@@ II Experimental results
For the three flow velocities, $U_{\text{c}}$, the three periods of boat motion, $T$, and for both types of propellers, rigid and folding, the angular velocity of the propeller, $\Omega$, required during the burst phase to keep the boat, on average, stationary in the flow increases as the duty cycle, $\Delta$, decreases (Fig. 3). The angular velocity is minimum for a duty cycle value of one, which corresponds to continuous rotation (no coast). As the duty cycle decreases, the propeller produces thrust for a shorter time, so the required angular velocity increases accordingly. Angular velocity shows little dependence on the period of motion compared to other parameters, i.e., flow velocity, duty cycle, or type of propeller. For a given duty cycle, the angular velocity increases with flow velocity; for continuous propulsion ($\Delta=1$), the tip speed ratio is constant ($R\Omega\propto U_{\text{c}}$, insert in Fig. 3), which we show theoretically in the Appendix. More importantly, the angular velocity of the folding propeller is always smaller than that of the rigid propeller. This difference increases as the duty cycle decreases, i.e., as the duration of the coast phase increases (relative to the period duration). This is due to the reconfiguration of the folding propeller, which reduces the drag force during the coast phase and the averaged energy cost for motion.
@@ III.1 Simple scaling
where $\eta$ accounts for the transfer from torque to thrust. The thrust works on a distance $\Delta U_{\rm{c}}$ and the drag force works over a distance $\Delta U_{\rm{c}}T$ with a coefficient $C_{D,\,\rm{B}}$ (burst, propeller in rotation) and over a distance $(1-\Delta)U_{\rm{c}}T$ with a coefficient $C_{D,\,\rm{C}}$ (coast, propeller at rest and folded when reconfigurable). This simple balance gives the following scaling for the angular velocity, $\Omega$, with duty cycle, $\Delta$, and drag contrast, $C_{D,\text{C}}/C_{D,\text{B}}$ :
@@ III.2 Full model
where $C_{a}$ is a function of $\lambda$: $C_{a}(\lambda)=a+c/\lambda^{2}$. The $a$ coefficient represents the thrust of the propeller and is positive, while the $c$ coefficient represents the quasi-steady drag of the boat and propeller and is negative ($C_{D,B}=-\pi R^{2}/S_{\text{boat}}c$). We have determined the values of these two coefficients experimentally in static conditions (See Supplementary Information).
@@
This model agrees well with experimental results for long period, as shown in Fig. 4 ($C_{\rm P}(\Delta)$) for $T=10\,s$. However, this model predicts that the gain increases (i.e. $C_{P}$ decreases), as the period (or $M$) decreases, contrary to what is observed in experiments. We believe this discrepancy is due to transient effects such as rotational inertia, which are not included in our model.
@@ IV Discussion
To discuss the efficiency gain of intermittent propulsion, we vary the main parameters of the system, the flow velocity and drag contrast, while keeping the period ($T=10\,\rm{s}$) and geometric parameters constant, and find the best theoretical relative efficiency gain:
@@
As shown in Figure 5, for all flow velocities (from 0 to 0.2 m/s), intermittent propulsion is more efficient than continuous propulsion when the drag contrast, $C_{D,\text{C}}/C_{D,\text{B}}$, is less than 0.35. Then the gain increases and the optimum duty cycle decreases as the drag contrast decreases. These theoretical results are confirmed by our experiments for two different drag contrasts. In the experiments, we vary the drag contrast by changing the relative importance of propeller and hull drag. The hull drag was increased by increasing the water height and the immersion of the hull. For a drag contrast of 0.2, the efficiency gain is between 15% and 24% for an optimum duty cycle of $0.35\pm 0.5$, depending on the flow velocity. When the drag contrast increases to 0.25, the optimum efficiency gain is only 7.5 % for a duty cycle of 0.55.
Figure 5: (a) Maximal relative mean power gains compared to continuous motion as a function of the drag contrast between the coast and burst phases. Error bars correspond to uncertainty in the data, which are estimated by repeating experiments. Inset: Comparison with fish swimming literature. (Blue) Pitching airfoil experiments from Floryan et al., 2017 [19] (Green) Experiments with carps from Wu et al., 2007 [4] (b) Duty cycle for which the power reduction is maximum. Error bars correspond to the range where the power coefficient is less than 5% away from the optimum. Data from experiments with water height ($\bullet$) 190 mm, ($\blacksquare$) 200 mm (See Fig. 4). (dashed line) Solution from Eq. 19. (full line) Solution with a constant torque coefficient (Eq. 20).
@@
as shown in the inset in Fig. 5. In particular, intermittent propulsion is favorable for a drag contrast less than $0.35$ compared to $2/3$ when the torque coefficient is independent of the tip speed ratio. This more favorable regime seems to be found in fish, as shown in the inset of Fig. 5, where we have plotted data from previous studies with live carp and pitching airfoil propulsion. Indeed, experimental data from Floryan et al. [24] show that the dimensionless cost of locomotion (equivalent to $C_{Q}$), when using a pitching airfoil, is independent of the actuation parameter (the Strouhal number, $St=fA/U_{\text{c}}$ where $f$ and $A$ are the tail-beat frequency and amplitude, respectively, analogous to the tip speed ratio). This particular dependence of the dimensionless cost of locomotion on the actuation parameter makes intermittent pitching particularly advantageous in fish propulsion. For a drag contrast of 0.25, the relative efficiency gain for carp is of 45% [4]. With a drag contrast of 0.55, intermittent propulsion is still beneficial ($\tau=0.1$) for pitching airfoils [19] (Fig. 5).
@@ V Conclusion
We have experimentally investigated the burst and coast dynamics in the case of a reconfigurable propeller, bio-inspired by fish swimming. Our work demonstrates that energy savings through intermittent motion can be achieved passively in artificial devices such as a propeller-driven boat. The conclusions of this work are applicable to any system driven through a fluid and subject to a typical high Reynolds number drag force that is quadratic in the speed of motion, such as fish or human swimming, bird flight, etc. Intermittent motion is advantageous when the drag contrast between the coast and burst phases is less than a threshold value, which is critically determined by the dependence of thrust cost on the actuation number, which compares the characteristic driving velocity to the moving speed, e.g. the tip speed number for propeller propulsion or the Strouhal number for fish swimming. For our propeller driven boat, for which cost of thrust increases with the tip speed ratio, intermittent motion is advantageous for drag contrasts less than 0.35. The onset increases to 2/3 when the thrust cost is independent of the actuation number, which is the case for pitching propulsion as for fish. A decreasing thrust cost with the actuation number would make burst and cost motion even more advantageous.
@@ Appendix A Materials and Methods
The boat is driven by a DC motor whose angular velocity is controlled to follow a prescribed time sequence. The torque is measured using a mechanical sensor mounted between the motor shaft and the propeller. Before each experiment, we measured the torque as a function of the angular velocity of the system without propeller, which we subtracted from the torque measured during the experiment (with the propeller) to only retain the hydrodynamic part. The boat is free to move along a linear air bearing rail (of negligible friction) in a 1.3 meter-long water channel of rectangular section 0.19× $0.19\,\times\,0.2\,{\rm m}$. The position of the boat is measured with an ultrasonic sensor. The water flow in the channel is generated by a pump of variable speed. The mean flow velocity, ranging from $0.05$ to $0.25$ m/s, is derived from measurements of water height and flux. A settling chamber, followed by a contraction and honeycomb flow straighteners dampen turbulent fluctuations at the channel entry.
@@
Figure 6: Three examples of boat trajectories in the laboratory reference frame. Colors represent the duty cycles (darker is longer). Here the propeller is rigid, $U_{\text{c}}=0.15$ m/s, and $T=5$ s. From $t=0$ the position of the boat is (quasi-) stationary on average.
@@
where $A$ and $B$ are integration constants, which are found numerically (root finding algorithm) to satisfy boundary conditions: the final velocity of one phase is the starting velocity of the other. For both types of propellers, we ran the simulations for the same parameters as in the experiments and for a period of 10 s (Fig. 4). As in the experiment, we looked for pairs $(\Delta,\Omega)$ such that the mean speed of the boat was equal to the prescribed value. Finally, we have verified that the main results presented in this manuscript are general and not strongly dependent on the propeller design. For a highly efficient marine propeller (parameters in Table 1 [25]), the model ((19)) shows that intermittent motion is advantageous for drag contrasts less than 0.3 (0.35 with the parameters of the reconfigurable propeller used in this study). For a drag contrast of 0.2, optimum power consumption is reduced by about 20% compared with continuous motion.

### v2

@@ Abstract
Many swimmers, especially small to medium-sized animals, use intermittent locomotion that differs from continuous swimming of large species. This type of locomotion, called burst-and-coast, is often associated with an energetic advantage. In this work, we investigate the intermittent locomotion inspired by fish locomotion but applied to a propeller. The energy consumption of burst-and-coast cycles is measured and compared to the continuous rotation regime. We show that a substantial drag ratio between the active and passive phases of the motion, as observed in fish, is critical for energy savings. Such a contrast can be obtained using a folding propeller that passively opens and closes as the propeller starts and stops rotating. For this reconfigurable propeller, intermittent propulsion is found to be energetically advantageous, saving up to 24% of the energy required to cruise at a given speed. Using an analytical model, we show that intermittent motion is more efficient than continuous motion when the drag reduction in the coast phase exceeds 65%. For fish-like locomotion, this threshold seems to be closer to 30%. A formal analogy allows us to explain the difference between propeller propulsion and fish locomotion.
@@ II Experimental results
For the three flow velocities, $U_{\text{c}}$, the three periods of boat motion, $T$, and for both types of propellers, rigid and folding, the angular velocity of the propeller, $\Omega$, required during the burst phase to keep the boat, on average, stationary in the flow increases as the duty cycle, $\Delta$, decreases (Fig. 3). The angular velocity is minimum for a duty cycle value of one, which corresponds to continuous rotation (no coast). As the duty cycle decreases, the propeller produces thrust for a shorter time, so the required angular velocity increases accordingly. Angular velocity shows little dependence on the period of motion compared to other parameters, i.e., flow velocity, duty cycle, or type of propeller. For a given duty cycle, the angular velocity increases with flow velocity; for continuous propulsion ($\Delta=1$), the tip speed ratio is constant ($R\Omega\propto U_{\text{c}}$, insert in Fig. 3), which we show theoretically in the Appendix. More importantly, the angular velocity of the folding propeller is always smaller than that of the rigid propeller. This difference increases as the duty cycle decreases, i.e., as the duration of the coast phase increases (relative to the period duration). This is due to the reconfiguration of the folding propeller, which reduces the drag force during the coast phase and the averaged energy cost for motion. More importantly, the angular velocity of the folding propeller is always smaller than that of the rigid propeller. This difference increases as the duty cycle decreases, i.e., as the duration of the coast phase increases (relative to the period duration). This is due to the reconfiguration of the folding propeller, which reduces the drag force during the coast phase and the averaged energy cost for motion.
@@
Insets show the same graph, but for the rigid propeller (Note that vertical scale is different).
@@ III.1 Simple scaling
where $\eta$ is the propulsive efficiency which takes into account the conversion from torque to thrust (the typical efficiency of our propellers is 0.3). The thrust works over a distance $\Delta U_{\rm{c}}T$ and the drag force works over a distance $\Delta U_{\rm{c}}T$ with a coefficient $C_{D,\,\rm{B}}$ (burst, propeller in rotation) and over a distance $(1-\Delta)U_{\rm{c}}T$ with a coefficient $C_{D,\,\rm{C}}$ (coast, propeller at rest and folded when reconfigurable). This simple balance gives the following scaling for the angular velocity, $\Omega$, with duty cycle, $\Delta$, and drag ratio, $C_{D,\text{C}}/C_{D,\text{B}}$ :
@@ III.2 Full model
where $C_{a}$ is a function of $\lambda$: $C_{a}(\lambda)=a+c/\lambda^{2}$. The $a$ coefficient represents the thrust of the propeller and is positive, while the $c$ coefficient represents the quasi-steady drag of the boat and propeller and is negative ($C_{D,B}=-\pi R^{2}/S_{\text{boat}}c$). We have determined the values of these two coefficients experimentally in static conditions and make the hypothesis that they remain unchanged during unsteady motion (See Supplementary Information).
@@
This model agrees well with experimental results for long period, as shown in Fig. 4 ($C_{\rm P}(\Delta)$) for $T=10\,s$. However, this model predicts that the gain increases (i.e. $C_{P}$ decreases), as the period (or $M$) decreases, contrary to what is observed in experiments. We believe this discrepancy is due to transient effects such as rotational inertia, which are not included in our model. Unsteady wave drag could also play a significant role. Depending on the period, wave drag has been shown to either increase or decrease the mean drag [11].
@@ IV Discussion
To discuss the efficiency gain of intermittent propulsion, we vary the main parameters of the system, the flow velocity and drag ratio, while keeping the period ($T=10\,\rm{s}$) and geometric parameters constant, and find the best theoretical relative efficiency gain:
@@
As shown in Figure 5, for all flow velocities (from 0 to 0.2 m/s), intermittent propulsion is more efficient than continuous propulsion when the drag ratio, $C_{D,\text{C}}/C_{D,\text{B}}$, is less than 0.35. Then the gain increases and the optimum duty cycle decreases as the drag ratio decreases. These theoretical results are confirmed by our experiments for two different drag ratios. In the experiments, we vary the drag ratio by changing the relative importance of propeller and hull drag. The hull drag was increased by increasing the water height and the immersion of the hull. For a drag ratio of 0.2, the efficiency gain is between 15% and 24% for an optimum duty cycle of $0.35\pm 0.5$, depending on the flow velocity. When the drag ratio increases to 0.25, the optimum efficiency gain is only 7.5 % for a duty cycle of 0.55.
Figure 5: (a) Maximal relative mean power gains compared to continuous motion as a function of the drag ratio between the coast and burst phases. Error bars correspond to uncertainty in the data, which are estimated by repeating experiments. Inset: Comparison with fish swimming literature. (Blue) Pitching airfoil experiments from Floryan et al., 2017 [19] (Green) Experiments with carps from Wu et al., 2007 [4] (b) Duty cycle for which the power reduction is maximum. Error bars correspond to the range where the power coefficient is less than 5% away from the optimum. Data from experiments with water height ($\bullet$) 190 mm, ($\blacksquare$) 200 mm (See Fig. 4). (dashed line) Solution from Eq. 19. (full line) Solution with a constant torque coefficient (Eq. 20).
@@
as shown in the inset in Fig. 5. In particular, intermittent propulsion is favorable for a drag ratio less than $0.35$ compared to $2/3$ when the torque coefficient is independent of the tip speed ratio. This more favorable regime seems to be found in fish, as shown in the inset of Fig. 5, where we have plotted data from previous studies with live carp [4] and pitching airfoil propulsion [24] ($Re\sim\mathcal{O}(10^{3}-10^{4})$ as in our study). Indeed, experimental data from Floryan et al. [24] show that the dimensionless cost of locomotion (equivalent to $C_{Q}$), when using a pitching airfoil, is independent of the actuation parameter (the Strouhal number, $St=fA/U_{\text{c}}$ where $f$ and $A$ are the tail-beat frequency and amplitude, respectively, analogous to the tip speed ratio). This particular dependence of the dimensionless cost of locomotion on the actuation parameter makes intermittent pitching particularly advantageous in fish propulsion. For a drag ratio of 0.25, the relative efficiency gain for carp is of 45% [4]. With a drag ratio of 0.55, intermittent propulsion is still beneficial ($\tau=0.1$) for pitching airfoils [19] (Fig. 5).
@@ V Conclusion
We have experimentally investigated the burst and coast dynamics in the case of a reconfigurable propeller, bio-inspired by fish swimming. Our work demonstrates that energy savings through intermittent motion can be achieved passively in artificial devices such as a propeller-driven boat. The conclusions of this work are applicable to any system driven through a fluid and subject to a typical high Reynolds number drag force that is quadratic in the speed of motion, such as fish or human swimming, bird flight, etc. Intermittent motion is advantageous when the drag ratio between the coast and burst phases is less than a threshold value, which is critically determined by the dependence of thrust cost on the actuation number, which compares the characteristic driving velocity to the moving speed, e.g. the tip speed number for propeller propulsion or the Strouhal number for fish swimming. For our propeller driven boat, for which cost of thrust increases with the tip speed ratio, intermittent motion is advantageous for drag ratios less than 0.35. The onset increases to 2/3 when the thrust cost is independent of the actuation number, which is the case for pitching propulsion as for fish. A decreasing thrust cost with the actuation number would make burst and cost motion even more advantageous.
@@ Appendix A Materials and Methods
The boat is driven by a DC motor whose angular velocity is controlled to follow a prescribed time sequence. The torque is measured using a mechanical sensor mounted between the motor shaft and the propeller. Before each experiment, we measured the torque as a function of the angular velocity of the system without propeller, which we subtracted from the torque measured during the experiment (with the propeller) to only retain the hydrodynamic part. The boat is free to move along a linear air bearing rail (of negligible friction) in a 1.3 meter-long water channel of rectangular section 0.19× $0.19\,\times\,0.2\,{\rm m}$. The position of the boat is measured with an ultrasonic sensor. The water flow in the channel is generated by a pump of variable speed. The mean flow velocity, ranging from $0.05$ to $0.25$ m/s, is derived from measurements of water height and flux. A settling chamber, followed by a contraction and honeycomb flow straighteners dampen turbulent fluctuations at the channel entry. The Reynolds number based on the size of the hull is about 15 000.
@@
Figure 6: Three examples of boat trajectories in the laboratory reference frame. Colors represent the duty cycles (darker is longer). Here the propeller is rigid, $U_{\text{c}}=0.15$ m/s, and $T=5$ s. From $t=0$ the position of the boat is (quasi-) stationary on average. (Inset) Root-mean-square fluctuations of the velocity as a function of $\mathcal{A}_{u}$. The green line is a fit with slope 0.18.
Given the intermittent nature of the motion, the velocity varies over time (Fig. 6). The amplitude of the velocity fluctuation $\mathcal{A}_{u}$, can be estimated using Eq. (10):
$\mathcal{A}_{u}=(1-\Delta)T\frac{1/2\rho S_{\textrm{boat}}C_{\textrm{D,C}}U_{c}^{2}}{m+m_{a}}.$ (21)
Eq. (21) fits well with the experimental data (inset of Fig. 6), collapsing data for different propellers, mean velocities, duty cycles, and periods.
@@
where $A$ and $B$ are integration constants, which are found numerically (root finding algorithm) to satisfy boundary conditions: the final velocity of one phase is the starting velocity of the other. For both types of propellers, we ran the simulations for the same parameters as in the experiments and for a period of 10 s (Fig. 4). As in the experiment, we looked for pairs $(\Delta,\Omega)$ such that the mean speed of the boat was equal to the prescribed value. Finally, we have verified that the main results presented in this manuscript are general and not strongly dependent on the propeller design. For a highly efficient marine propeller (parameters in Table 1 [25]), the model ((19)) shows that intermittent motion is advantageous for drag ratios less than 0.3 (0.35 with the parameters of the reconfigurable propeller used in this study). For a drag ratio of 0.2, optimum power consumption is reduced by about 20% compared with continuous motion.